A 5'6" person walking down the street notices his shadow. If the angle of elevation from the tip of their shadow to the sun is 60°, what is the distance from the tip of the shadow to the top of his head (round to 2 decimal places)?
Answer by ewatrrr(24785)

 
Hi,
A 5'6" person walking down the street notices his shadow.
If the angle of elevation from the tip of their shadow to the sun is 60°
what is the distancex from the tip of the shadow to the top of his head
sin 60 = opp/hyp = 66in/x
sin 60 = 66/x
x = 66/sin 60 = 66/.866 = 76.21 in